### v1

# The Spectral Flow of a Restriction to a Subspace

Abstract. We prove in full generality a formula that relates the spectral flow of a continuous path of self-adjoint Fredholm operators with the spectral flow of the restrictions of the operators to a fixed closed finite codimensional subspace.
1.    Introduction
Given a real finite dimensional vector space $E$, a quadratic form $Q$ on $E$ and a vector subspace $W\subset E$, the index of $Q$ differs from the index of its restriction to $W$ by the following expression
${\rm ind}\,Q-{\rm ind}\,Q|_{W}={\rm ind}\,Q|_{W^{\perp_{Q}}}+\dim(W\cap W^{\perp_{Q}})-\dim(W\cap\ker Q),$ (1)
where $W^{\perp_{Q}}=\left\{X\in E\,\;|\,\;Q[X,Y]=0\;\;\forall Y\in W\right\}$. This linear algebra fact played a key role in the establishing in [2] of a Morse index theorem for closed Riemannian geodesics from the theorem for geodesic segments.
When dealing with strongly indefinite problems (e.g. when working on a semi-Riemannian background), the notion of Morse index is replaced by the notion of spectral flow (see Sec. 2.2), and it becomes relevant to extend formula (1) to a formula relating spectral flows. For instance, in order to work out a Morse index theorem for closed semi-Riemannian geodesics, the authors of [3] established a formula (see Eq. (2) below) relating the spectral flow of a path $\{Q_{t}\}_{t\in[a,b]}$ of quadratic forms of Fredholm type on a Hilbert space $H$ to the spectral flow of the path of the restrictions $Q_{t}|_{V}$ to a fixed closed finite codimensional subspace $V\subset H$. The same formula was also derived in [6] in order to establish a Morse index theorem for electromagnetic geodesic segments constrained to a fixed energy level. While the proof in [3] works only for paths $\{Q_{t}\}$ whose corresponding path $\{L_{t}\}$ of self-adjoint operators is a perturbation of a fixed symmetry $\mathfrak{I}\in GL(H)$ by a path of compact operators (in which case the spectral flow of $\{Q_{t}\}$ depends only on the endpoints $Q_{a}$ and $Q_{b}$), the proof in [6] works without this restriction but assumes non-degeneracy of the paths $\{Q_{t}\}$ and $\{Q_{t}|_{V}\}$ at the endpoints. Here we shall extend the proof in [6] so as to allow for degenerate endpoints. More precisely, we shall prove the following:
Theorem 1.1.
Let $Q=\{Q_{t}\}_{t\in[a,b]}$ be a continuous path of quadratic forms of Fredholm type on a separable infinite dimensional real Hilbert space $H$. Let $V\subset H$ be a closed finite codimensional subspace and denote by $\overline{Q}=\{\overline{Q}_{t}\}_{t\in[a,b]}$ the path of quadratic forms on $V$ obtained by restricting the $Q_{t}$ to $V$. Then the spectral flows ${\rm sf}(Q,[a,b])$ and ${\rm sf}(\overline{Q},[a,b])$ are related by
${\rm sf}(Q,[a,b])-{\rm sf}(\overline{Q},[a,b])$ $\;=\;{\rm ind}\,Q_{a}|_{V^{\perp_{Q_{a}}}}-{\rm ind}\,Q_{b}|_{V^{\perp_{Q_{b}}}}+{\rm dim}(V\cap V^{\perp_{Q_{a}}})-{\rm dim}(V\cap V^{\perp_{Q_{b}}})$ (2)
$\;\quad-{\rm dim}(V\cap\ker Q_{a})+{\rm dim}(V\cap\ker Q_{b}).$
In Sec. 2 we collect some properties of the spectral flow (Sec. 2.2) and prove some preliminary results on quadratic forms of Fredholm type (Sec. 2.1) that shall be used for the proof of Theorem 1.1; in particular, we prove a result on the genericity of the non-degenerate subspaces with respect to such a quadratic form. Sec. 3 contains the proof of Theorem 1.1.
2.    Preliminaries
Throughout these notes, $H$ shall denote a separable infinite dimensional real Hilbert space. A bounded quadratic form $Q$ on $H$ is of Fredholm type if the bounded self-adjoint operator that represents it is Fredholm. It is clear that the restriction of such a form to a closed finite codimensional subspace of $H$ is also of Fredholm type.
Recall that a quadratic form $Q$ is said non-degenerate if $\ker Q=H\cap H^{\perp_{Q}}=\{0\}$, and that a subspace $W\subset H$ is non-degenerate with respect to $Q$ if $\ker Q|_{W}=W\cap W^{\perp_{Q}}=\{0\}$. If $Q$ is of Fredholm type and $W\subset H$ is finite codimensional (respectively, finite dimensional), then $W^{\perp_{Q}}$ is finite dimensional (respectively, finite codimensional) as one easily shows.
2.1    Some results on quadratic forms.
Let $Q$ be a fixed quadratic form of Fredholm type on $H$.
Lemma 2.1.
Let $U\subset H$ be a subspace that is either finite dimensional or closed and finite codimensional. Then, $(U^{\perp_{Q}})^{\perp_{Q}}=U+\ker Q$.
Proof. Let $\pi:H\rightarrow H/\ker Q$ be the quotient map and $\overline{Q}$ be the quadratic form on $H/\ker Q$ induced by $Q$. Then,
  1. 1.
    $\bigl{(}W^{\perp_{\overline{Q}}}\bigr{)}^{\perp_{\overline{Q}}}=W$ for any closed subspace $W\subset H/\ker Q$ since the musical map $\overline{Q}^{\sharp}:H/\ker Q\rightarrow(H/\ker Q)^{*}$ is an isomorphism;
  2. 2.
    $\pi(V^{\perp_{Q}})=\pi(V)^{\perp_{\overline{Q}}}$ for any subspace $V\subset H$, by a purely algebraic argument;
  3. 3.
    $\pi(U)$ is closed since the restriction $\pi|_{U}:U\rightarrow H/\ker Q$ is a Fredholm operator if $U$ is closed and ${\rm codim}\,U<\infty$ (since $\pi$ is already Fredholm) and Fredholm operators have closed range.
Therefore we can apply 1. to $W=\pi(U)$ and iterate 2. to obtain $\pi\bigl{(}(U^{\perp_{Q}})^{\perp_{Q}}\bigr{)}=\pi(U)$. From this equality and the fact that $\ker Q\subset(U^{\perp_{Q}})^{\perp_{Q}}$, we conclude the result.
Lemma 2.2.
Let $W\subset V$ be closed subspaces of $H$ with finite codimension and with $W$ non-degenerate. Then, $(W^{\perp_{Q}}\cap V)^{\perp_{Q}}=W\oplus V^{\perp_{Q}}$.
Proof. Fist of all, observe that $E=W\oplus V^{\perp_{Q}}$ is closed in $H$: indeed, if $\pi:H\rightarrow W/W$ is the quotient map, then $\pi$ is continuous, $\pi(E)$ is finite dimensional (hence, closed) and $W=\pi^{-1}(\pi(E))$. Now, it is clear that $(W\oplus V^{\perp_{Q}})^{\perp_{Q}}=W^{\perp_{Q}}\cap(V^{\perp_{Q}})^{\perp_{Q}}$. Thus, since, by the previous lemma, $(V^{\perp_{Q}})^{\perp_{Q}}=V+\ker Q$, and since $\ker Q\subset W^{\perp_{Q}}$, we obtain $(W\oplus V^{\perp_{Q}})^{\perp_{Q}}=W^{\perp_{Q}}\cap V+\ker Q$. On the other hand, since $W^{\perp_{Q}}\cap V$ is finite dimensional, Lemma 2.1 also applies to show that $\left((W^{\perp_{Q}}\cap V)^{\perp_{Q}}\right)^{\perp_{Q}}=W^{\perp_{Q}}\cap V+\ker Q$. Therefore, $\left((W^{\perp_{Q}}\cap V)^{\perp_{Q}}\right)^{\perp_{Q}}=(W\oplus V^{\perp_{Q}})^{\perp_{Q}}$. Computing the $\perp_{Q}$ of both sides of last equality via Lemma 2.1, we obtain $(W^{\perp_{Q}}\cap V)^{\perp_{Q}}+\ker Q=W\oplus V^{\perp_{Q}}+\ker Q$. The result now follows since $\ker Q$ is contained in both $W^{\perp_{Q}}\cap V$ and $W\oplus V^{\perp_{Q}}$.
For the proof of Theorem 1.1 we shall need to guarantee the existence of a closed finite codimensional subspace that is non-degenerate with respect to two quadratic forms (see Lemma 3.3 of Sec. 3). The easiest way to do so seems to be establishing a much stronger result about the distribution of non-degenerate subspaces:
Lemma 2.3.
The non-degenerate closed subspaces $W\subset H$ of a fixed finite codimension $k$ form an open subset of the Grassmann manifold $G^{k}(H)$ of all $k-$codimensional closed subspaces of $H$. If $Q$ is non-degenerate, then that subset is also dense.
Proof. We shall also denote by $G_{k}(H)$ the Grassmannian of all $k-$dimensional subspaces of $H$. Let us first prove the assertion about density. This is equivalent to proving the density in $G_{k}(H)$ of the subset $\Sigma_{k}$ of non-degenerate subspaces. For, the non-degeneracy of $Q$ implies that a subspace $W\in G^{k}(H)$ is non-degenerate if, and only if, $W^{\perp_{Q}}$ is non-degenerate (this follows from Lemma 2.1) and that the map $G^{k}(H)\rightarrow G_{k}(H)$, $W\mapsto W^{\perp_{Q}}$, is a homeomorphism. The proof of the density of $\Sigma_{k}$ in $G_{k}(H)$ is an application of the infinite dimensional Transversality Density Theorem (TDT) ([1, Thm. 19.1]) as we now explain. Let $\mathscr{C}=\{X\in H\backslash\{0\}\;\,|\;\,Q[X]=0\}$ be the isotropic cone of $Q$. Then a subspace $W\subset H$ is degenerate if, and only if, $W$ is somewhere tangent to $\mathscr{C}$, i.e. there exists $X\in W\cap\mathscr{C}$ such that $W$ is contained in the tangent space $T_{X}\mathscr{C}$ (this follows since $T_{X}\mathscr{C}=\langle X\rangle^{\perp_{Q}}$). Passing to the projective space $\mathds{P}(H)$, which is a smooth Hilbert manifold, that condition is equivalent to the (finite dimensional) projective subspace $\mathds{P}(W)\subset\mathds{P}(H)$ being somewhere tangent to the projective “hiperquadric” $\mathds{P}(\mathscr{C})\subset\mathds{P}(H)$, which is a one codimensional smooth submanifold of $\mathds{P}(H)$. In order to apply TDT, we need to put the set of embeddings $\left\{\mathds{P}(W)\hookrightarrow\mathds{P}(H)\;\,|\,\;W\in G_{k}(H)\right\}$ in a parametric form (at least in a neighborhood of each point of $G_{k}(H)$). For this, given $W_{0}\in G_{k}(H)$, let $L(W_{0},W_{0}^{\perp})$ be the Hilbertable space of all (necessarily bounded) linear maps $W_{0}\rightarrow W_{0}^{\perp}$. The set of $W\in G_{k}(H)$ with $W\cap W_{0}^{\perp}=\{0\}$ forms a neighborhood of $W_{0}$ on which a smooth parametrization $L(W_{0},W_{0}^{\perp})\rightarrow{\cal U}_{W_{0}}$ is given by $f\mapsto{\rm graph}\,f=\{X+f(X)\,\;|\,\;X\in W_{0}\}$. We then define
$F:L(W_{0},W_{0}^{\perp})\times\mathds{P}(W_{0})\rightarrow\mathds{P}(H),\quad F(f,[X])=[X+f(X)],$
where []delimited-[][-][ - ] stands for the projective class of a vector. The map $F$ is transversal to $\mathds{P}(\mathscr{C})$ since it is a submersion, hence TDT applies to guarantee that the set of $f$ for which $\mathds{P}({\rm graph}\,f)\hookrightarrow\mathds{P}(H)$ is transversal to $\mathds{P}(\mathscr{C})$ is dense. Therefore, the set of $W\in G_{k}(H)$ for which $\mathds{P}(W)\hookrightarrow\mathds{P}(H)$ is transverse to $\mathds{P}(\mathscr{C})$ is dense.
Let us now consider the assertion about openness. Given $W_{0}\in G^{k}(H)$ with $W_{0}\cap W_{0}^{\perp_{Q}}=\{0\}$, there exist neighborhoods ${\cal A}$ of $W_{0}$ in $G_{k}(H)$ and ${\cal A}^{\prime}$ of $W_{0}^{\perp_{Q}}$ in $G^{k}(H)$ such that $W\cap W^{\prime}=\{0\}$ for all $W\in{\cal A}$ and $W^{\prime}\in{\cal A}^{\prime}$. On the other hand, since $Q$ may be degenerate, the correspondence $W\mapsto W^{\perp_{Q}}$ is not continuous on the whole of $G_{k}(H)$ (indeed, $\dim W^{\perp_{Q}}$ might jump). Nevertheless, one can show without difficulty that on the neighborhood ${\cal V}$ of $W_{0}$ given by ${\cal V}=\{W\,\;|\,\;W\cap\ker Q=\{0\}\}$ the map $W\mapsto W^{\perp_{Q}}$ takes values in $G_{k}(H)$ and is continuous. Thus, for $W$ sufficiently close to $W_{0}$ we will have $W^{\perp_{Q}}\in{\cal A}^{\prime}$. This concludes the proof.
2.2    The spectral flow.
Let $\hat{\mathfrak{F}}=\hat{\mathfrak{F}}(H)$ be the space of self-adjoint Fredholm operators on $H$. The space $\hat{\mathfrak{F}}$ has three connected components: two contractible components $\hat{\mathfrak{F}}_{+}$ and $\hat{\mathfrak{F}}_{-}$, which consist on the operators in $\hat{\mathfrak{F}}$ that are, respectively, essentially positive and essentially negative (i.e. that have finite dimensional negative, respectively, positive, spectral subspaces), and one component $\hat{\mathfrak{F}}_{*}$ that has fundamental group isomorphic to $\mathds{Z}$; operators in $\hat{\mathfrak{F}}_{*}$ are called strongly indefinite. Let now $L=\{L_{t}\}_{t\in[a,b]}$ be a continuous path in $\hat{\mathfrak{F}}$; equivalently, we can consider the corresponding path $Q=\{Q_{t}\}_{t\in[a,b]}$ of quadratic forms of Fredholm type on $H$. The spectral flow of $L$, or of $Q$, is the integer ${\rm sf}(L,[a,b])$, or ${\rm sf}(Q,[a,b])$, which, intuitively, gives the net number of eigenvalues of $L$ that cross 0 in the positive direction as $t$ runs from $a$ to $b$. An intuitive, albeit rigorous, construction of ${\rm sf}(L,[a,b])$ that works regardless of endpoints invertibility assumptions was carried out by Phillips in [5] by means of the continuous functional calculus. According to this definition, which we shall assume, if $L$ is a path in $\hat{\mathfrak{F}}_{+}$ then
${\rm sf}(L,[a,b])={\rm ind}\,L_{a}-{\rm ind}\,L_{b}.$ (3)
We shall not need the precise definition of ${\rm sf}(L,[a,b])$ in what follows, but just some basic properties that we collect bellow and that can be found in [4].
  1. 1.
    If $L_{t}$ is invertible for all $t$, then ${\rm sf}(L,[a,b])=0$;
  2. 2.
    if $L_{1}*L_{2}$ is the concatenation of two paths $L_{1}=\{(L_{1})_{t}\}_{t\in[a,b]}$ and $L_{2}=\{(L_{2})_{t}\}_{t\in[b,c]}$ that agree for $t=b$, then
    ${\rm sf}(L_{1}*L_{2},[a,c])={\rm sf}(L_{1},[a,b])+{\rm sf}(L_{2},[b,c]);$
  3. 3.
    if $L_{1}$ and $L_{2}$ are two paths of self-adjoint Fredholm operators on Hilbert spaces $H_{1}$ and $H_{2}$, then
    ${\rm sf}(L_{1}\oplus L_{2},[a,b])={\rm sf}(L_{1},[a,b])+{\rm sf}(L_{2},[a,b]);$
  4. 4.
    given a path $L$ in $\hat{\mathfrak{F}}(H)$ and a path $M=\{M_{t}\}$ in $GL(H)$, then
    ${\rm sf}(L,[a,b])={\rm sf}(M^{*}LM,[a,b]);$
  5. 5.
    if $L$ is closed and $K=\{K_{t}\}_{t\in[a,b]}$ is a closed path of compact self-adjoint operators, then
    ${\rm sf}(L,[a,b])={\rm sf}(L+K,[a,b]).$
Observe that, from properties 1., 3. and 4, we also have
Lemma 2.4.
If the kernel of $L_{t}$ has constant dimension, then ${\rm sf}(L,[a,b])=0$.
Proof. Let $H_{1}=\ker L_{a}$. Since $\dim\ker L_{t}$ is constant, it is easy to show that we can find a continuous path $M=\{M_{t}\}$ in $GL(H)$ such that $M_{t}(H_{1})=\ker L_{t}$ for all $t$. The operator $M_{t}^{*}L_{t}M_{t}$ then splits as a direct sum $0\oplus L_{t}^{\prime}$, where 00 is the null operator on $H_{1}$ end $L_{t}^{\prime}$ is an invertible operator on $H_{1}^{\perp}$. The result follows now from 1., 3. and 4.
3.    Proof of Theorem 1.1
We shall only deal with the relevant case in which $Q$ is a path of strongly indefinite quadratic forms. In this case, it is clear that $\{\overline{Q}_{t}\}$ are also strongly indefinite. The important fact in this regard will be the fact, used in the proof of Lemma 3.2 below, that the space $\hat{\mathfrak{F}}_{*}(V)\cap GL(V)$ is path-connected (see [6, Lemma 3.2]).
As in [6], the proof of Theorem 1.1 is based on the following observation, whose simple proof we shall reproduce here for the convenience of the reader:
Lemma 3.1.
Suppose that the path $Q$ is closed. Then ${\rm sf}(Q,[a,b])={\rm sf}(\overline{Q},[a,b])$.
Proof. Let $L_{t}$ be the self-adjoint operator representing $Q_{t}$. We split $L_{t}$ as $L_{t}=N_{t}+K_{t}$, where the operators are given in block forms, relative to the decomposition $H=V\oplus V^{\perp}$, as
$N_{t}=\begin{pmatrix}A_{t}&{\rm O}\\ {\rm O}&{\rm O}\end{pmatrix},\quad K_{t}=\begin{pmatrix}{\rm O}&B_{t}\\ B_{t}^{*}&C_{t}\end{pmatrix}.$
The self-adjoint representation of $\overline{Q}_{t}$ is $A_{t}$, and since $V^{\perp}$ is finite dimensional, $K_{t}$ is compact. It thus follows from 3. and 5. of Sec. 2.2 that ${\rm sf}(N,[a,b])={\rm sf}(\overline{Q},[a,b])$ and ${\rm sf}(N+K,[a,b])={\rm sf}(N,[a,b])$.
Lemma 3.2.
If $\overline{Q}_{a}$ and $\overline{Q}_{b}$ are non-degenerate, then
${\rm sf}(Q,[a,b])-{\rm sf}(\overline{Q},[a,b])={\rm ind}\,Q_{a}|_{V^{\perp_{Q_{a}}}}-{\rm ind}\,Q_{b}|_{V^{\perp_{Q_{b}}}}.$
Proof. The proof is essentially the same as that of [6, Prop. 3.4], except that this time we need Lemma 2.4. The idea is to close up the path $Q=\{Q_{t}\}_{t\in[a,b]}$ as follows. Since $\overline{Q}_{a}$ and $\overline{Q}_{b}$ are non-degenerate, we have $H=V\oplus V^{\perp_{Q_{a}}}=V\oplus V^{\perp_{Q_{b}}}$. So, for $s=a,b$, by rotating $V^{\perp_{Q_{s}}}$ to $V^{\perp}$ while keeping $V$ fixed, we can connect $Q_{s}$, via a path of quadratic forms of Fredholm type, to the quadratic form $\hat{Q}_{s}$ whose block decomposition relative to $H=V\oplus V^{\perp}$ is
$\hat{Q}_{s}=\begin{pmatrix}\overline{Q}_{s}&{\rm O}\\ {\rm O}&Q_{s}^{\perp}\end{pmatrix},\quad s=a,b,$
for some quadratic form $Q_{s}^{\perp}$ on $V^{\perp}$ which is isometric to $Q_{s}|_{V^{\perp_{Q_{s}}}}$. Both paths have constant dimensional kernels, so by Lemma 2.4 their spectral flows are equal to zero. Next we connect $\hat{Q}_{b}$ to $\hat{Q}_{a}$ by, at the same time, connecting $\overline{Q}_{b}$ to $\overline{Q}_{a}$ via a path of non-degenerate quadratic forms on $V$ (this is possible since the space $\hat{\mathfrak{F}}_{*}(V)\cap GL(V)$ is path-connected) and connecting, in any way, $Q_{b}^{\perp}$ to $Q_{a}^{\perp}$ and keeping the null blocks unaltered. By additivity under direct sum, the spectral flow of this path is the sum of the spectral flows of its restrictions to $V$ and $V^{\perp}$, respectively. The first restriction is a path of non-degenerate quadratic forms, so has null spectral flow. The second is a path of quadratic forms on the finite dimensional Hilbert space $V^{\perp}$ starting at $Q^{\perp}_{b}$ and ending at $Q^{\perp}_{a}$, hence its spectral flow is simply given by ${\rm ind}\,Q^{\perp}_{b}-{\rm ind}\,Q^{\perp}_{a}={\rm ind}\,Q_{b}|_{V^{\perp_{Q_{b}}}}-{\rm ind}\,Q_{a}|_{V^{\perp_{Q_{a}}}}$ (see Eq. (3)). Now, by concatenating to $Q$ the above paths we obtain a closed path $Q^{\prime}=\{Q^{\prime}_{t}\}_{t\in[a,c]}$, for some $b<c$, such that $Q^{\prime}_{t}=Q_{t}$ for $t\in[a,b]$. From the above discussion, and from the additivity of the spectral flow under concatenation, we have
${\rm sf}(Q^{\prime},[a,c])={\rm sf}(Q,[a,b])+{\rm ind}\,Q_{b}|_{V^{\perp_{Q_{b}}}}-{\rm ind}\,Q_{a}|_{V^{\perp_{Q_{a}}}}.$
On the other hand, since $Q^{\prime}$ is a closed path, Lemma 3.1 implies that ${\rm sf}(Q^{\prime},[a,c])={\rm sf}(\overline{Q^{\prime}},[a,c])$, where $\overline{Q^{\prime}}$ is the restriction of $Q^{\prime}$ to $V$. But since $\overline{Q^{\prime}}_{t}$ is non-degenerate for $t\in[b,c]$, then ${\rm sf}(\overline{Q^{\prime}},[a,c])={\rm sf}(\overline{Q^{\prime}},[a,b])={\rm sf}(\overline{Q},[a,b])$. The result follows.
Lemma 3.3.
There exists a closed finite codimensional subspace $W\subset H$ that is contained in $V$ and which is non-degenerate with respect to both $Q_{a}$ and $Q_{b}$.
Proof. Let $W^{\prime}\subset V$ be any closed subspace such that $V=W^{\prime}\oplus\ker\overline{Q}_{a}$. Then $Q_{a}$ is non-degenerate on $W^{\prime}$ and $W^{\prime}$ has finite codimension in $H$. By the same reasoning, we can find a closed finite codimensional subspace $W^{\prime\prime}\subset H$ with $W^{\prime\prime}\subset W^{\prime}$ and on which $Q_{b}$ is non-degenerate. Let $k$ be the codimension of $W^{\prime\prime}$ in $W^{\prime}$. From Lemma 2.3, there exists a neighborhood ${\cal U}$ of $W^{\prime\prime}$ in $G^{k}(W^{\prime})$ such that $Q_{b}$ is non-degenerate on any $W\in{\cal U}$. Also, since $Q_{a}$ is non-degenerate on $W^{\prime}$, the second part of Lemma 2.3 applies to guarantee a $W\in{\cal U}$ on which $Q_{a}$ is non-degenerate. This completes the proof.
Let $W$ be as in the previous lemma. We can apply Lemma 3.2 twice: one time with $V$ replaced by $W$, and the other with $H$ replaced by $V$ (and $Q$ replaced by $Q|_{V}$) and $V$ replaced by $W$. We obtain, respectively,
[…]
Therefore,
${\rm sf}(Q,[a,b])-{\rm sf}(\overline{Q},[a,b])=\left({\rm ind}\,Q_{a}|_{W^{\perp_{Q_{a}}}}-{\rm ind}\,Q_{a}|_{W^{\perp_{Q_{a}}}\cap V}\right)-\left({\rm ind}\,Q_{b}|_{W^{\perp_{Q_{b}}}}-{\rm ind}\,Q_{b}|_{W^{\perp_{Q_{b}}}\cap V}\right).$ (4)
From now on, let $s=a,b$. Each difference of indices above can be computed by means of Eq. (1) since $W^{\perp_{Q_{s}}}$ is finite dimensional. Doing so, we obtain
${\rm ind}\,Q_{s}|_{W^{\perp_{Q_{s}}}}-{\rm ind}\,Q_{s}|_{W^{\perp_{Q_{s}}}\cap V}$ $\;=\;{\rm ind}\left(Q_{s}\big{|}_{(W^{\perp_{Q_{s}}}\cap V)^{\perp_{Q_{s}}}\cap W^{\perp_{Q_{s}}}}\right)+\dim\left((W^{\perp_{Q_{s}}}\cap V)\cap(W^{\perp_{Q_{s}}}\cap V)^{\perp_{Q_{s}}}\right)$ (5)
$\quad-\dim\left((W^{\perp_{Q_{s}}}\cap V)\cap\ker Q_{s}|_{W^{\perp_{Q_{s}}}}\right).$
The intersections of subspaces in the right-hand side above can be simplified as follows. On the one hand, since $H=W\oplus W^{\perp_{Q_{s}}}$ and $\ker Q_{s}\subset W^{\perp_{Q_{s}}}$, we have $\ker Q_{s}|_{W^{\perp_{Q_{s}}}}=\ker Q_{s}$ and thus $(W^{\perp_{Q_{s}}}\cap V)\cap\ker Q_{s}|_{W^{\perp_{Q_{s}}}}=V\cap\ker Q_{s}$. On the other hand, from Lemma 2.2 we obtain
$(W^{\perp_{Q_{s}}}\cap V)\cap(W^{\perp_{Q_{s}}}\cap V)^{\perp_{Q_{s}}}$ $=$ $(W^{\perp_{Q_{s}}}\cap V)\cap(W\oplus V^{\perp_{Q_{s}}})$
$=$ $V\cap V^{\perp_{Q_{s}}},$
$(W^{\perp_{Q_{s}}}\cap V)^{\perp_{Q_{s}}}\cap W^{\perp_{Q_{s}}}$ $=$ $(W\oplus V^{\perp_{Q_{s}}})\cap W^{\perp_{Q_{s}}}$
$=$ $V^{\perp_{Q}}.$
The second and fourth equalities above are direct consequences of the relations $W\subset V$ and $W^{\perp_{Q_{s}}}\supset V^{\perp_{Q_{s}}}$ and $W\cap W^{\perp_{Q_{s}}}=\{0\}$. Eq. (5) can thus be rewritten as
${\rm ind}\,Q_{s}|_{W^{\perp_{Q_{s}}}}-{\rm ind}\,Q_{s}|_{W^{\perp_{Q_{s}}}\cap V}={\rm ind}\,Q_{s}|_{V^{\perp_{Q_{s}}}}+\dim(V\cap V^{\perp_{Q_{s}}})-\dim(V\cap\ker Q_{s}).$
Substituting the above, for $s=a$ and $s=b$, in the right-hand side of Eq. (4) concludes the proof of Theorem 1.1.
Data availability statement
No new data were created or analyzed in this study. Data sharing is not applicable to this article.