### v2

# The Spectral Flow of a Restriction to a Subspace and

Abstract. We prove in full generality a formula that relates the spectral flow of a continuous path of quadratic forms of Fredholm type with the spectral flow of the restrictions of the forms to a fixed closed finite codimensional subspace. We then apply this to obtain a formula relating the Maslov index of a continuous path in a Fredholm Lagrangian Grassmannian with the Maslov index of its symplectic reduction by a closed finite codimensional coisotropic subspace.
1.    Introduction
Given a real finite dimensional vector space $E$, a quadratic form $Q$ on $E$ and a vector subspace $W\subset E$, the index of $Q$ differs from the index of its restriction to $W$ by the following expression
${\rm ind}\,Q-{\rm ind}\,Q|_{W}={\rm ind}\,Q|_{W^{\perp_{Q}}}+\dim(W\cap W^{\perp_{Q}})-\dim(W\cap\ker Q),$ (1)
where $W^{\perp_{Q}}=\left\{X\in E\,\;|\,\;Q[X,Y]=0\;\;\forall Y\in W\right\}$. This linear algebra fact played a key role in the establishing in [2] of a Morse index theorem for closed Riemannian geodesics from the theorem for geodesic segments.
When dealing with strongly indefinite problems (e.g. when working on a semi-Riemannian background), the notion of Morse index is replaced by the notion of spectral flow (see Sec. 3), and it becomes relevant to extend formula (1) to a formula relating spectral flows. In these notes, we shall prove the following
Theorem 1.1.
Let $Q=\{Q_{t}\}_{t\in[a,b]}$ be a continuous path of quadratic forms of Fredholm type on a separable infinite dimensional real Hilbert space $H$. Let $V\subset H$ be a closed finite codimensional subspace and denote by $\overline{Q}=\{\overline{Q}_{t}\}$ the path of quadratic forms on $V$ obtained by restricting the $Q_{t}$ to $V$. Then the spectral flows ${\rm sf}(Q,[a,b])$ and ${\rm sf}(\overline{Q},[a,b])$ are related by
${\rm sf}(Q,[a,b])-{\rm sf}(\overline{Q},[a,b])$ $\;=\;{\rm ind}\,Q_{a}|_{V^{\perp_{Q_{a}}}}-{\rm ind}\,Q_{b}|_{V^{\perp_{Q_{b}}}}+{\rm dim}(V\cap V^{\perp_{Q_{a}}})-{\rm dim}(V\cap V^{\perp_{Q_{b}}})$ (2)
$\;\quad-\dim(V\cap\ker Q_{a})+\dim(V\cap\ker Q_{b}).$
It turns out that being able to relate the spectral flows of $\{Q_{t}\}$ and $\{\overline{Q}_{t}\}$ plays a key role in many places. For instance,
  • 1–
    Under the hypothesis that the path $\{L_{t}\}$ of self-adjoint operators representing $\{Q_{t}\}$ is a perturbation of a fixed symmetry $\mathfrak{I}\in GL(H)$ by a path of compact operators (in which case the spectral flow of $\{Q_{t}\}$ depends only on the endpoints $Q_{a}$ and $Q_{b}$), formula (2) was proved in [3] and used to work out a Morse index theorem for closed semi-Riemannian geodesics;
  • 2–
    For establishing in [10] a Morse index theorem for electromagnetic geodesic segments constrained to a fixed energy level, one had to resort to such a formula. The proof of formula (2) given in [10] worked without the restriction in [3] but assumed non-degeneracy of the paths $\{Q_{t}\}$ and $\{\overline{Q}_{t}\}$ at the endpoints;
  • 3–
    As we shall show in these notes, when performing a symplectic reduction of a path $\ell:[a,b]\rightarrow{\cal F}\Lambda_{L_{0}}$ in a Fredholm Lagrangian Grassmannian (see Sec. 4), formula (2) leads naturally to a formula relating the Maslov index $\mu_{L_{0}}(\ell,[a,b])$ of $\ell$ with the Maslov index of the reduction of $\ell$.
The proper context of 3– is the following. Let $\mathds{H}$ be a symplectic Hilbertable space and let $\mathds{W}\subset\mathds{H}$ be a closed finite codimensional coisotropic subspace. The quotient $\mathds{W}/\mathds{W}^{\perp_{\omega}}$ then becomes a Hilbertable symplectic space. Let $q:\mathds{W}\rightarrow\mathds{W}/\mathds{W}^{\perp_{\omega}}$ be the quotient map. Then we shall prove:
Theorem 1.2.
Let $L_{0}\in\Lambda(\mathds{H})$ have null intersection with $\mathds{W}^{\perp_{\omega}}$ and let $\ell:[a,b]\rightarrow{\cal F}\Lambda_{L_{0}}(\mathds{H})$ be a continuous path with $\ell(t)\cap\mathds{W}^{\perp_{\omega}}=\{0\}$ for all $t$. Set $\overline{L}_{0}=q(L_{0}\cap\mathds{W})$. Then $t\mapsto\overline{\ell}(t)=q(\ell(t)\cap\mathds{W})$ defines a continuous path in ${\cal F}\Lambda_{\overline{L}_{0}}(\mathds{W}/\mathds{W}^{\perp_{\omega}})$ and
$\mu_{L_{0}}(\ell,[a,b])-\mu_{\overline{L}_{0}}(\overline{\ell},[a,b])$ $\;=\;{\rm ind}\,\mathfrak{q}_{a}-{\rm ind}\,\mathfrak{q}_{b}+\dim\pi\bigl{(}\ell(a)\cap(V+\mathds{W}^{\perp_{\omega}})\bigr{)}-\dim\pi\bigl{(}\ell(b)\cap(V+\mathds{W}^{\perp_{\omega}})\bigr{)}$
$\;\quad-\dim(\ell(a)\cap V)+\dim(\ell(b)\cap V).$
In the above,
  1. 1.
    $V=L_{0}\cap\mathds{W}$, so $V$ is a closed finite codimensional subspace of $L_{0}$;
  2. 2.
    $\pi:L_{0}+\mathds{W}^{\perp_{\omega}}\rightarrow L_{0}$ is the projection;
  3. 3.
    for $t=a,b$, $\mathfrak{q}_{t}$ is the quadratic form on the finite dimensional space $E_{t}=\pi\bigl{(}\ell(t)\cap(L_{0}+\mathds{W}^{\perp_{\omega}})\bigr{)}$ given by
    $\mathfrak{q}_{t}[u,v]=\omega\Bigl{[}\left(\pi\big{|}_{E_{t}}\right)^{-1}u,v\Bigr{]}.$
The above result should be compared with [8, Prop. 1.15] in which an equality of Maslov indices is obtained in a suitable symplectic reduction.
This work is organized as follows. In Sec. 2 we recall some concepts and prove some results about bilinear forms (symmetric and skew-symmetric) of Fredholm type. After collecting some properties of the spectral flow in Sec. 3, we prove Theorem 1.1 in Sec. 3.1. In Sec. 4 we begin by recalling the concepts surrounding the Fredholm Lagrangian Grassmannian and the Maslov index of paths. The proof of Theorem 1.2 is then carried out in Sec. 4.2 and Sec. 4.3.
2.    Some preliminaries
Throughout this section, $H$ shall denote a separable infinite dimensional real Hilbertable space. Let $Q$ be a bounded bilinear form on $H$ which is either symmetric (in which case we shall call it a quadratic form) or skew-symmetric. For a subspace $W\subseteq H$, we set $W^{\perp_{Q}}=\{u\in H\,\;|\,\;Q[u,w]=0\;\;\forall w\in W\}$. Recall that $Q$ is said non-degenerate if $\ker Q=H\cap H^{\perp_{Q}}=\{0\}$, and that a subspace $W\subset H$ is non-degenerate with respect to $Q$ if $\ker Q|_{W}=W\cap W^{\perp_{Q}}=\{0\}$. The form $Q$ is said of Fredholm type if the bounded self-adjoint (or skew-adjoint) operator that represents it with respect to some (and hence to all) compatible Hilbert inner product on $H$ is Fredholm. It is clear that if $Q$ is of Fredholm type and if $W\subset H$ is a closed finite codimensional subspace, then the restriction $Q|_{W}$ is of Fredholm type and $W^{\perp_{Q}}$ is finite dimensional. Also, if $Q$ is non-degenerate and of Fredholm type, then $Q$ is strongly non-degenerate, i.e. the musical map $Q^{\sharp}:H\rightarrow H^{*}$ is an isomorphism.
For the next two lemmas, let $Q$ be a fixed quadratic/skew-symmetric form of Fredholm type on $H$.
Lemma 2.1.
Let $U\subset H$ be a subspace that is either finite dimensional or closed and finite codimensional. Then, $(U^{\perp_{Q}})^{\perp_{Q}}=U+\ker Q$.
Proof. Let $\pi:H\rightarrow H/\ker Q$ be the quotient map and $\overline{Q}$ be the form on $H/\ker Q$ induced by $Q$. Then,
  1. 1.
    $\bigl{(}W^{\perp_{\overline{Q}}}\bigr{)}^{\perp_{\overline{Q}}}=W$ for any closed subspace $W\subset H/\ker Q$ since the musical map $\overline{Q}^{\sharp}:H/\ker Q\rightarrow(H/\ker Q)^{*}$ is an isomorphism;
  2. 2.
    $\pi(V^{\perp_{Q}})=\pi(V)^{\perp_{\overline{Q}}}$ for any subspace $V\subset H$, by a purely algebraic argument;
  3. 3.
    $\pi(U)$ is closed since the restriction $\pi|_{U}:U\rightarrow H/\ker Q$ is a Fredholm operator if $U$ is closed and ${\rm codim}\,U<\infty$ (since $\pi$ is already Fredholm) and Fredholm operators have closed range.
Therefore we can apply 1. to $W=\pi(U)$ and iterate 2. to obtain $\pi\bigl{(}(U^{\perp_{Q}})^{\perp_{Q}}\bigr{)}=\pi(U)$. From this equality and the fact that $\ker Q\subset(U^{\perp_{Q}})^{\perp_{Q}}$, we conclude the result.
Lemma 2.2.
Let $W\subset V$ be closed subspaces of $H$ with finite codimension and with $W$ non-degenerate. Then, $(W^{\perp_{Q}}\cap V)^{\perp_{Q}}=W\oplus V^{\perp_{Q}}$.
Proof. Fist of all, observe that $E=W\oplus V^{\perp_{Q}}$ is closed in $H$: indeed, if $\pi:H\rightarrow H/W$ is the quotient map, then $\pi$ is continuous, $\pi(E)$ is finite dimensional (hence, closed) and $W=\pi^{-1}(\pi(E))$. Now, it is clear that $(W\oplus V^{\perp_{Q}})^{\perp_{Q}}=W^{\perp_{Q}}\cap(V^{\perp_{Q}})^{\perp_{Q}}$. Thus, since, by the previous lemma, $(V^{\perp_{Q}})^{\perp_{Q}}=V+\ker Q$, and since $\ker Q\subset W^{\perp_{Q}}$, we obtain $(W\oplus V^{\perp_{Q}})^{\perp_{Q}}=W^{\perp_{Q}}\cap V+\ker Q$. On the other hand, since $W^{\perp_{Q}}\cap V$ is finite dimensional, Lemma 2.1 also applies to show that $\left((W^{\perp_{Q}}\cap V)^{\perp_{Q}}\right)^{\perp_{Q}}=W^{\perp_{Q}}\cap V+\ker Q$. Therefore, $\left((W^{\perp_{Q}}\cap V)^{\perp_{Q}}\right)^{\perp_{Q}}=(W\oplus V^{\perp_{Q}})^{\perp_{Q}}$. Computing the $\perp_{Q}$ of both sides of last equality via Lemma 2.1, we obtain $(W^{\perp_{Q}}\cap V)^{\perp_{Q}}+\ker Q=W\oplus V^{\perp_{Q}}+\ker Q$. The result now follows since $\ker Q$ is contained in both $W^{\perp_{Q}}\cap V$ and $W\oplus V^{\perp_{Q}}$.
For the proof of Theorem 1.1 we shall need to guarantee the existence of a closed finite codimensional subspace that is non-degenerate with respect to two quadratic forms (see Lemma 3.4 of Sec. 3.1). The easiest way to do so seems to be establishing a much stronger result about the distribution of non-degenerate subspaces:
Lemma 2.3.
Let be fixed a quadratic form $Q$ of Fredholm type on $H$. The non-degenerate closed subspaces $W\subset H$ of a fixed finite codimension $k$ form an open subset of the Grassmann manifold $G^{k}(H)$ of all $k-$codimensional closed subspaces of $H$. If $Q$ is non-degenerate, then that subset is also dense.
Proof. We shall also denote by $G_{k}(H)$ the Grassmannian of all $k-$dimensional subspaces of $H$. Let us first prove the assertion about density. This is equivalent to proving the density in $G_{k}(H)$ of the subset $\Sigma_{k}$ of non-degenerate subspaces. For, the non-degeneracy of $Q$ implies that a subspace $W\in G^{k}(H)$ is non-degenerate if, and only if, $W^{\perp_{Q}}$ is non-degenerate (this follows from Lemma 2.1) and that the map $G^{k}(H)\rightarrow G_{k}(H)$, $W\mapsto W^{\perp_{Q}}$, is a homeomorphism. The proof of the density of $\Sigma_{k}$ in $G_{k}(H)$ is an application of the infinite dimensional Transversality Density Theorem (TDT) ([1, Thm. 19.1]) as we now explain. Let $\mathscr{C}=\{X\in H\backslash\{0\}\;\,|\;\,Q[X]=0\}$ be the isotropic cone of $Q$. Then a subspace $W\subset H$ is degenerate if, and only if, $W$ is somewhere tangent to $\mathscr{C}$, i.e. there exists $X\in W\cap\mathscr{C}$ such that $W$ is contained in the tangent space $T_{X}\mathscr{C}$ (this follows since $T_{X}\mathscr{C}=\langle X\rangle^{\perp_{Q}}$). Passing to the projective space $\mathds{P}(H)$, which is a smooth Hilbert manifold, that condition is equivalent to the (finite dimensional) projective subspace $\mathds{P}(W)\subset\mathds{P}(H)$ being somewhere tangent to the projective “hiperquadric” $\mathds{P}(\mathscr{C})\subset\mathds{P}(H)$, which is a one codimensional smooth submanifold of $\mathds{P}(H)$. In order to apply TDT, we need to put the set of embeddings $\left\{\mathds{P}(W)\hookrightarrow\mathds{P}(H)\;\,|\,\;W\in G_{k}(H)\right\}$ in a parametric form (at least in a neighborhood of each point of $G_{k}(H)$). For this, given $W_{0}\in G_{k}(H)$, let $L(W_{0},W_{0}^{\perp})$ be the Hilbertable space of all (necessarily bounded) linear maps $W_{0}\rightarrow W_{0}^{\perp}$. The set of $W\in G_{k}(H)$ with $W\cap W_{0}^{\perp}=\{0\}$ forms a neighborhood of $W_{0}$ on which a smooth parametrization $L(W_{0},W_{0}^{\perp})\rightarrow{\cal U}_{W_{0}}$ is given by $f\mapsto{\rm graph}\,f=\{X+f(X)\,\;|\,\;X\in W_{0}\}$. We then define
$F:L(W_{0},W_{0}^{\perp})\times\mathds{P}(W_{0})\rightarrow\mathds{P}(H),\quad F(f,[X])=[X+f(X)],$
where []delimited-[][-][ - ] stands for the projective class of a vector. The map $F$ is transversal to $\mathds{P}(\mathscr{C})$ since it is a submersion, hence TDT applies to guarantee that the set of $f$ for which $\mathds{P}({\rm graph}\,f)\hookrightarrow\mathds{P}(H)$ is transversal to $\mathds{P}(\mathscr{C})$ is dense. Therefore, the set of $W\in G_{k}(H)$ for which $\mathds{P}(W)\hookrightarrow\mathds{P}(H)$ is transverse to $\mathds{P}(\mathscr{C})$ is dense.
Let us now consider the assertion about openness. Given $W_{0}\in G^{k}(H)$ with $W_{0}\cap W_{0}^{\perp_{Q}}=\{0\}$, there exist neighborhoods ${\cal A}$ of $W_{0}$ in $G_{k}(H)$ and ${\cal A}^{\prime}$ of $W_{0}^{\perp_{Q}}$ in $G^{k}(H)$ such that $W\cap W^{\prime}=\{0\}$ for all $W\in{\cal A}$ and $W^{\prime}\in{\cal A}^{\prime}$. On the other hand, since $Q$ may be degenerate, the correspondence $W\mapsto W^{\perp_{Q}}$ is not continuous on the whole of $G_{k}(H)$ (indeed, $\dim W^{\perp_{Q}}$ might jump). Nevertheless, one can show without difficulty that on the neighborhood ${\cal V}$ of $W_{0}$ given by ${\cal V}=\{W\,\;|\,\;W\cap\ker Q=\{0\}\}$ the map $W\mapsto W^{\perp_{Q}}$ takes values in $G_{k}(H)$ and is continuous. Thus, for $W$ sufficiently close to $W_{0}$ we will have $W^{\perp_{Q}}\in{\cal A}^{\prime}$. This concludes the proof.
3.    The spectral flow
Let $H$ be a separable infinite dimensional real Hilbert space, and let $\hat{\mathfrak{F}}=\hat{\mathfrak{F}}(H)$ be the space of self-adjoint Fredholm operators on $H$. The space $\hat{\mathfrak{F}}$ has three connected components: two contractible components $\hat{\mathfrak{F}}_{+}$ and $\hat{\mathfrak{F}}_{-}$, which consist on the operators in $\hat{\mathfrak{F}}$ that are, respectively, essentially positive and essentially negative (i.e. that have finite dimensional negative, respectively, positive, spectral subspaces), and one component $\hat{\mathfrak{F}}_{*}$ that has fundamental group isomorphic to $\mathds{Z}$; operators in $\hat{\mathfrak{F}}_{*}$ are called strongly indefinite. Let now $L=\{L_{t}\}_{t\in[a,b]}$ be a continuous path in $\hat{\mathfrak{F}}$; equivalently, we can consider the corresponding path $Q=\{Q_{t}\}_{t\in[a,b]}$ of quadratic forms of Fredholm type on $H$. The spectral flow of $L$, or of $Q$, is the integer ${\rm sf}(L,[a,b])$, or ${\rm sf}(Q,[a,b])$, which, intuitively, gives the net number of eigenvalues of $L$ that cross 0 in the positive direction as $t$ runs from $a$ to $b$. An intuitive, albeit rigorous, construction of ${\rm sf}(L,[a,b])$ that works regardless of endpoints invertibility assumptions was carried out by Phillips in [9] by means of the continuous functional calculus. According to this definition, which we shall assume, if $L$ is a path in $\hat{\mathfrak{F}}_{+}$ then
${\rm sf}(L,[a,b])={\rm ind}\,L_{a}-{\rm ind}\,L_{b}.$ (3)
We shall not need the precise definition of ${\rm sf}(L,[a,b])$ in what follows, but just some basic properties that we collect bellow and that can be found in [6].
  1. 1.
    If $L_{t}$ is invertible for all $t$, then ${\rm sf}(L,[a,b])=0$;
  2. 2.
    if $L_{1}*L_{2}$ is the concatenation of two paths $L_{1}=\{(L_{1})_{t}\}_{t\in[a,b]}$ and $L_{2}=\{(L_{2})_{t}\}_{t\in[b,c]}$ that agree for $t=b$, then
    ${\rm sf}(L_{1}*L_{2},[a,c])={\rm sf}(L_{1},[a,b])+{\rm sf}(L_{2},[b,c]);$
  3. 3.
    if $L_{1}$ and $L_{2}$ are two paths of self-adjoint Fredholm operators on Hilbert spaces $H_{1}$ and $H_{2}$, then
    ${\rm sf}(L_{1}\oplus L_{2},[a,b])={\rm sf}(L_{1},[a,b])+{\rm sf}(L_{2},[a,b]);$
  4. 4.
    given a path $L$ in $\hat{\mathfrak{F}}(H)$ and a path $M=\{M_{t}\}$ in $GL(H)$, then
    ${\rm sf}(L,[a,b])={\rm sf}(M^{*}LM,[a,b]);$
  5. 5.
    if $L$ is closed and $K=\{K_{t}\}_{t\in[a,b]}$ is a closed path of compact self-adjoint operators, then
    ${\rm sf}(L,[a,b])={\rm sf}(L+K,[a,b]).$
Observe that, from properties 1., 3. and 4, we also have
Lemma 3.1.
If the kernel of $L_{t}$ has constant dimension, then ${\rm sf}(L,[a,b])=0$.
Proof. Let $H_{1}=\ker L_{a}$. Since $\dim\ker L_{t}$ is constant, it is easy to show that we can find a continuous path $M=\{M_{t}\}$ in $GL(H)$ such that $M_{t}(H_{1})=\ker L_{t}$ for all $t$. The operator $M_{t}^{*}L_{t}M_{t}$ then splits as a direct sum $0\oplus L_{t}^{\prime}$, where 00 is the null operator on $H_{1}$ end $L_{t}^{\prime}$ is an invertible operator on $H_{1}^{\perp}$. The result follows now from 1., 3. and 4.
3.1    Proof of Theorem 1.1
We shall only deal with the relevant case in which $Q$ is a path of strongly indefinite quadratic forms. In this case, it is clear that $\{\overline{Q}_{t}\}$ are also strongly indefinite. The important fact in this regard will be the fact, used in the proof of Lemma 3.3 below, that the space $\hat{\mathfrak{F}}_{*}(V)\cap GL(V)$ is path-connected (see [10, Lemma 3.2]).
As in [10], the proof of Theorem 1.1 is based on the following observation, whose simple proof we shall reproduce here for the convenience of the reader:
Lemma 3.2.
Suppose that the path $Q$ is closed. Then ${\rm sf}(Q,[a,b])={\rm sf}(\overline{Q},[a,b])$.
Proof. Let $L_{t}$ be the self-adjoint operator representing $Q_{t}$. We split $L_{t}$ as $L_{t}=N_{t}+K_{t}$, where the operators are given in block forms, relative to the decomposition $H=V\oplus V^{\perp}$, as
$N_{t}=\begin{pmatrix}A_{t}&{\rm O}\\ {\rm O}&{\rm O}\end{pmatrix},\quad K_{t}=\begin{pmatrix}{\rm O}&B_{t}\\ B_{t}^{*}&C_{t}\end{pmatrix}.$
The self-adjoint representation of $\overline{Q}_{t}$ is $A_{t}$, and since $V^{\perp}$ is finite dimensional, $K_{t}$ is compact. It thus follows from 3. and 5. of Sec. 3 that ${\rm sf}(N,[a,b])={\rm sf}(\overline{Q},[a,b])$ and ${\rm sf}(N+K,[a,b])={\rm sf}(N,[a,b])$.
Lemma 3.3.
If $\overline{Q}_{a}$ and $\overline{Q}_{b}$ are non-degenerate, then
${\rm sf}(Q,[a,b])-{\rm sf}(\overline{Q},[a,b])={\rm ind}\,Q_{a}|_{V^{\perp_{Q_{a}}}}-{\rm ind}\,Q_{b}|_{V^{\perp_{Q_{b}}}}.$
Proof. The proof is essentially the same as that of [10, Prop. 3.4], except that this time we need Lemma 3.1. The idea is to close up the path $Q=\{Q_{t}\}_{t\in[a,b]}$ as follows. Since $\overline{Q}_{a}$ and $\overline{Q}_{b}$ are non-degenerate, we have $H=V\oplus V^{\perp_{Q_{a}}}=V\oplus V^{\perp_{Q_{b}}}$. So, for $s=a,b$, by rotating $V^{\perp_{Q_{s}}}$ to $V^{\perp}$ while keeping $V$ fixed, we can connect $Q_{s}$, via a path of quadratic forms of Fredholm type, to the quadratic form $\hat{Q}_{s}$ whose block decomposition relative to $H=V\oplus V^{\perp}$ is
$\hat{Q}_{s}=\begin{pmatrix}\overline{Q}_{s}&{\rm O}\\ {\rm O}&Q_{s}^{\perp}\end{pmatrix},\quad s=a,b,$
for some quadratic form $Q_{s}^{\perp}$ on $V^{\perp}$ which is isometric to $Q_{s}|_{V^{\perp_{Q_{s}}}}$. Both paths have constant dimensional kernels, so by Lemma 3.1 their spectral flows are equal to zero. Next we connect $\hat{Q}_{b}$ to $\hat{Q}_{a}$ by, at the same time, connecting $\overline{Q}_{b}$ to $\overline{Q}_{a}$ via a path of non-degenerate quadratic forms on $V$ (this is possible since the space $\hat{\mathfrak{F}}_{*}(V)\cap GL(V)$ is path-connected) and connecting, in any way, $Q_{b}^{\perp}$ to $Q_{a}^{\perp}$ and keeping the null blocks unaltered. By additivity under direct sum, the spectral flow of this path is the sum of the spectral flows of its restrictions to $V$ and $V^{\perp}$, respectively. The first restriction is a path of non-degenerate quadratic forms, so has null spectral flow. The second is a path of quadratic forms on the finite dimensional Hilbert space $V^{\perp}$ starting at $Q^{\perp}_{b}$ and ending at $Q^{\perp}_{a}$, hence its spectral flow is simply given by ${\rm ind}\,Q^{\perp}_{b}-{\rm ind}\,Q^{\perp}_{a}={\rm ind}\,Q_{b}|_{V^{\perp_{Q_{b}}}}-{\rm ind}\,Q_{a}|_{V^{\perp_{Q_{a}}}}$ (see Eq. (3)). Now, by concatenating to $Q$ the above paths we obtain a closed path $Q^{\prime}=\{Q^{\prime}_{t}\}_{t\in[a,c]}$, for some $b<c$, such that $Q^{\prime}_{t}=Q_{t}$ for $t\in[a,b]$. From the above discussion, and from the additivity of the spectral flow under concatenation, we have
${\rm sf}(Q^{\prime},[a,c])={\rm sf}(Q,[a,b])+{\rm ind}\,Q_{b}|_{V^{\perp_{Q_{b}}}}-{\rm ind}\,Q_{a}|_{V^{\perp_{Q_{a}}}}.$
On the other hand, since $Q^{\prime}$ is a closed path, Lemma 3.2 implies that ${\rm sf}(Q^{\prime},[a,c])={\rm sf}(\overline{Q^{\prime}},[a,c])$, where $\overline{Q^{\prime}}$ is the restriction of $Q^{\prime}$ to $V$. But since $\overline{Q^{\prime}}_{t}$ is non-degenerate for $t\in[b,c]$, then ${\rm sf}(\overline{Q^{\prime}},[a,c])={\rm sf}(\overline{Q^{\prime}},[a,b])={\rm sf}(\overline{Q},[a,b])$. The result follows.
Lemma 3.4.
There exists a closed finite codimensional subspace $W\subset H$ that is contained in $V$ and which is non-degenerate with respect to both $Q_{a}$ and $Q_{b}$.
Proof. Let $W^{\prime}\subset V$ be any closed subspace such that $V=W^{\prime}\oplus\ker\overline{Q}_{a}$. Then $Q_{a}$ is non-degenerate on $W^{\prime}$ and $W^{\prime}$ has finite codimension in $H$. By the same reasoning, we can find a closed finite codimensional subspace $W^{\prime\prime}\subset H$ with $W^{\prime\prime}\subset W^{\prime}$ and on which $Q_{b}$ is non-degenerate. Let $k$ be the codimension of $W^{\prime\prime}$ in $W^{\prime}$. From Lemma 2.3, there exists a neighborhood ${\cal U}$ of $W^{\prime\prime}$ in $G^{k}(W^{\prime})$ such that $Q_{b}$ is non-degenerate on any $W\in{\cal U}$. Also, since $Q_{a}$ is non-degenerate on $W^{\prime}$, the second part of Lemma 2.3 applies to guarantee a $W\in{\cal U}$ on which $Q_{a}$ is non-degenerate. This completes the proof.
Let $W$ be as in the previous lemma. We can apply Lemma 3.3 twice: one time with $V$ replaced by $W$, and the other with $H$ replaced by $V$ (and $Q$ replaced by $Q|_{V}$) and $V$ replaced by $W$. We obtain, respectively,
[…]
Therefore,
${\rm sf}(Q,[a,b])-{\rm sf}(\overline{Q},[a,b])=\left({\rm ind}\,Q_{a}|_{W^{\perp_{Q_{a}}}}-{\rm ind}\,Q_{a}|_{W^{\perp_{Q_{a}}}\cap V}\right)-\left({\rm ind}\,Q_{b}|_{W^{\perp_{Q_{b}}}}-{\rm ind}\,Q_{b}|_{W^{\perp_{Q_{b}}}\cap V}\right).$ (4)
From now on, let $s=a,b$. Each difference of indices above can be computed by means of Eq. (1) since $W^{\perp_{Q_{s}}}$ is finite dimensional. Doing so, we obtain
${\rm ind}\,Q_{s}|_{W^{\perp_{Q_{s}}}}-{\rm ind}\,Q_{s}|_{W^{\perp_{Q_{s}}}\cap V}$ $\;=\;{\rm ind}\left(Q_{s}\big{|}_{(W^{\perp_{Q_{s}}}\cap V)^{\perp_{Q_{s}}}\cap W^{\perp_{Q_{s}}}}\right)+\dim\left((W^{\perp_{Q_{s}}}\cap V)\cap(W^{\perp_{Q_{s}}}\cap V)^{\perp_{Q_{s}}}\right)$ (5)
$\quad-\dim\left((W^{\perp_{Q_{s}}}\cap V)\cap\ker Q_{s}|_{W^{\perp_{Q_{s}}}}\right).$
The intersections of subspaces in the right-hand side above can be simplified as follows. On the one hand, since $H=W\oplus W^{\perp_{Q_{s}}}$ and $\ker Q_{s}\subset W^{\perp_{Q_{s}}}$, we have $\ker Q_{s}|_{W^{\perp_{Q_{s}}}}=\ker Q_{s}$ and thus $(W^{\perp_{Q_{s}}}\cap V)\cap\ker Q_{s}|_{W^{\perp_{Q_{s}}}}=V\cap\ker Q_{s}$. On the other hand, from Lemma 2.2 we obtain
$(W^{\perp_{Q_{s}}}\cap V)\cap(W^{\perp_{Q_{s}}}\cap V)^{\perp_{Q_{s}}}$ $=$ $(W^{\perp_{Q_{s}}}\cap V)\cap(W\oplus V^{\perp_{Q_{s}}})$
$=$ $V\cap V^{\perp_{Q_{s}}},$
$(W^{\perp_{Q_{s}}}\cap V)^{\perp_{Q_{s}}}\cap W^{\perp_{Q_{s}}}$ $=$ $(W\oplus V^{\perp_{Q_{s}}})\cap W^{\perp_{Q_{s}}}$
$=$ $V^{\perp_{Q}}.$
The second and fourth equalities above are direct consequences of the relations $W\subset V$ and $W^{\perp_{Q_{s}}}\supset V^{\perp_{Q_{s}}}$ and $W\cap W^{\perp_{Q_{s}}}=\{0\}$. Eq. (5) can thus be rewritten as
${\rm ind}\,Q_{s}|_{W^{\perp_{Q_{s}}}}-{\rm ind}\,Q_{s}|_{W^{\perp_{Q_{s}}}\cap V}={\rm ind}\,Q_{s}|_{V^{\perp_{Q_{s}}}}+\dim(V\cap V^{\perp_{Q_{s}}})-\dim(V\cap\ker Q_{s}).$
Substituting the above, for $s=a$ and $s=b$, in the right-hand side of Eq. (4) concludes the proof of Theorem 1.1.
4.    The Maslov index in a symplectic reduction
Throughout this section, $\mathds{H}$ shall denote a symplectic Hilbertable space. By this we mean that $\mathds{H}$ is a separable infinite dimensional real Hilbertable space endowed with a bounded skew-symmetric bilinear form $\omega$ which is strongly non-degenerate.
4.1    The Fredholm Lagrangian Grassmannian and the Maslov index.
Here we shall recall some notions regarding the Fredholm Lagrangian Grassmannian, ${\cal F}\Lambda_{L_{0}}(\mathds{H})$, and state a definition of the Maslov index $\mu_{L_{0}}(\ell,[a,b])$ for continous paths $\ell:[a,b]\rightarrow{\cal F}\Lambda_{L_{0}}(\mathds{H})$ in terms of the fundamental groupoid as presented in [4].
A (necessarily closed) subspace $L\subseteq\mathds{H}$ is said to be Lagrangian if $L=L^{\perp_{\omega}}$. The collection of all Lagrangian subspaces of $\mathds{H}$, the Lagrangian Grassmannian of $\mathds{H}$, is denoted by $\Lambda=\Lambda(\mathds{H})$. An analytic Banach manifold structure for $\Lambda$ is obtained as follows. Firstly, given $L\in\Lambda$, let $\Lambda_{0}(L)$ be the set of all Lagrangians $L^{\prime}$ that are complementary to $L$, i.e. $L^{\prime}\cap L=\{0\}$ and $L+L^{\prime}=\mathds{H}$; this is an open and dense subset of $\Lambda$. Given a pair $(L_{0},L_{1})$ of complementary Lagrangians, every $L\in\Lambda_{0}(L_{1})$ is the graph of a unique bounded linear map $T:L_{0}\rightarrow L_{1}$ and the bounded bilinear form on $L_{0}$ defined by $\omega[T\cdot,\cdot]|_{L_{0}\times L_{0}}$ is symmetric. This defines a chart for $\Lambda$ around $L_{0}$,
[…]
where $B_{\rm sym}(L_{0})$ denotes the (Banach) space of all bounded symmetric bilinear forms on $L_{0}$.
The Fredholm Lagrangian Grassmannian corresponding to a given $L_{0}\in\Lambda$ is the subset ${\cal F}\Lambda_{L_{0}}\subset\Lambda$ given by the $L$ such that $(L_{0},L)$ is a Fredholm pair, i.e. such that $L_{0}\cap L$ is finite dimensional and $L_{0}+L$ is closed and finite codimensional. According to [5, Lemma 2.6], one has
Lemma 4.1.
Given a Lagrangian $L_{1}$ complementary to $L_{0}$, a Lagrangian $L\in\Lambda_{0}(L_{1})$ forms a Fredholm pair with $L_{0}$ if, and only if, the quadratic form $\varphi_{L_{0},L_{1}}(L)$ is of Fredholm type.
It follows that ${\cal F}\Lambda_{L_{0}}$ is an open subset of $\Lambda$. While the full Lagrangian Grassmannian $\Lambda$ is contractible (due to Kuiper’s theorem), its Fredholm counterpart ${\cal F}\Lambda_{L_{0}}$ has an infinite cyclic fundamental group (see [8]). An isomorphism $\pi_{1}({\cal F}\Lambda_{L_{0}})\cong\mathds{Z}$ is provided by the Maslov index, whose definition can be given via the following result from [4] (see also [5]).
Theorem 4.2.
For each $L_{0}\in\Lambda$, there exists a unique function assigning to each continuous path $\ell:[a,b]\rightarrow{\cal F}\Lambda_{L_{0}}$ an integer $\mu_{L_{0}}(\ell,[a,b])$ that satisfies
  1. 1.
    $\mu_{L_{0}}$ is fixed-endpoint homotopy invariant;
  2. 2.
    $\mu_{L_{0}}$ is additive by concatenation;
  3. 3.
    if $\ell([a,b])\subset\Lambda_{0}(L_{1})$ for some $L_{1}$ complementary to $L_{0}$, and if $Q=\{Q_{t}\}$ is the path of quadratic forms of Fredholm type on $L_{0}$ given by $Q_{t}=\varphi_{L_{0},L_{1}}(\ell(t))$, then
    $\mu_{L_{0}}(\ell,[a,b])={\rm sf}(Q,[a,b]).$
4.2    Symplectic reduction.
Let $\mathds{W}\subset\mathds{H}$ be a closed finite codimensional subspace which is coisotropic, i.e. $\mathds{W}\supseteq\mathds{W}^{\perp_{\omega}}$. The restriction of $\omega$ to $\mathds{W}$ then descends to a bounded skew-symmetric and non-degenerate form $\overline{\omega}$ on $\mathds{W}/\mathds{W}^{\perp_{\omega}}$ because $\ker\omega|_{\mathds{W}\times\mathds{W}}=\mathds{W}^{\perp_{\omega}}$. Since ${\rm codim}\,\mathds{W}<\infty$, the form $\omega|_{\mathds{W}\times\mathds{W}}$ is of Fredholm type. Thus $\overline{\omega}$ is also of Fredholm type and, therefore, it is strongly non-degenerate. The pair $(\mathds{W}/\mathds{W}^{\perp_{\omega}},\overline{\omega})$ is called the symplectic reduction of $\mathds{H}$ by $\mathds{W}$.
We shall be interested in the Lagrangians $L\in\Lambda(\mathds{H})$ having null intersection with $\mathds{W}^{\perp_{\omega}}$. So let ${\cal O}\subset\Lambda(\mathds{H})$ be the open set given by
${\cal O}=\{L\in\Lambda(\mathds{H})\,\;|\,\;L\cap\mathds{W}^{\perp_{\omega}}=\{0\}\}.$
Let $q:\mathds{W}\rightarrow\mathds{W}/\mathds{W}^{\perp_{\omega}}$ be the quotient map.
Lemma 4.3.
We have a well-defined continuous map $\lambda:{\cal O}\rightarrow\Lambda(\mathds{W}/\mathds{W}^{\perp_{\omega}})$, $\lambda(L)=q(L\cap\mathds{W})$. Moreover, for a fixed $L_{0}\in{\cal O}$, we have $q(L\cap\mathds{W})\in{\cal F}\Lambda_{q(L_{0}\cap\mathds{W})}$ if $L\in{\cal F}\Lambda_{L_{0}}(\mathds{H})$.
Proof. Given $L_{0}\in{\cal O}$ we can find $L_{1}\in\Lambda_{0}(L_{0})$ such that $L_{1}\supset\mathds{W}^{\perp_{\omega}}$, and hence $L_{1}\subset\mathds{W}$ (because $L_{1}^{\perp_{\omega}}=L_{1}$ and $(\mathds{W}^{\perp_{\omega}})^{\perp_{\omega}}=\mathds{W}$). Set $V=\mathds{W}\cap L_{0}$. Since $\mathds{H}=L_{0}\oplus L_{1}$ and $L_{1}\subset\mathds{W}$, then $\mathds{W}=V\oplus L_{1}$. It follows that $\mathds{W}/\mathds{W}^{\perp_{\omega}}$ is canonically $V\oplus(L_{1}/\mathds{W}^{\perp_{\omega}})$, with symplectic form $\overline{\omega}\bigl{[}(u,v+\mathds{W}^{\perp_{\omega}}),(u^{\prime},v^{\prime}+\mathds{W}^{\perp_{\omega}})\bigr{]}=\omega[u,v^{\prime}]+\omega[v,u^{\prime}]$ and with $\lambda(L_{0})$ being the summand $V$. In particular, it is clear that $\lambda(L_{0})$ is Lagrangian. Set $\overline{L}_{0}=V$ and $\overline{L}_{1}=L_{1}/\mathds{W}^{\perp_{\omega}}$. Then $\mathds{W}/\mathds{W}^{\perp_{\omega}}=\overline{L}_{0}\oplus\overline{L}_{1}$ is a Lagrangian decomposition.
Claim. We have $\lambda(\Lambda_{0}(L_{1}))\subset\Lambda_{0}(\overline{L}_{1})$, and the map $\lambda:\Lambda_{0}(L_{1})\rightarrow\Lambda_{0}(\overline{L}_{1})$ corresponds, via the charts $\varphi_{L_{0},L_{1}}$ and $\varphi_{\overline{L}_{0},\overline{L}_{1}}$, to the map $B_{\rm sym}(L_{0})\rightarrow B_{\rm sym}(\overline{L}_{0})$, $Q\mapsto Q|_{V=\overline{L}_{0}}$.
Proof. Given $L\in\Lambda_{0}(L_{1})$, let $T:L_{0}\rightarrow L_{1}$ be such that $L=\{u+Tu\,\;|\,\;u\in L_{0}\}$, so that $\varphi_{L_{0},L_{1}}(L)$ is the quadratic form $Q[u,v]=\omega[Tu,v]$. From $L_{1}\subset\mathds{W}$ we obtain $L\cap\mathds{W}=\{u+Tu\,\;|\,\;u\in L_{0}\cap\mathds{W}=W\}$. It follows that $q(L\cap\mathds{W})$ is the graph of $\overline{L}_{0}\rightarrow\overline{L}_{1}$, $u\mapsto Tu+\mathds{W}^{\perp_{\omega}}$. Therefore, $\lambda(L)\in\Lambda_{0}(\overline{L}_{1})$ and $\varphi_{\overline{L}_{0},\overline{L}_{1}}(\lambda(L))$ is the quadratic form $\overline{Q}$ on $V=\overline{L}_{0}$ given by $\overline{Q}[u,v]=\overline{\omega}\bigl{[}(0,Tu+\mathds{W}^{\perp_{\omega}}),(v,0)\bigr{]}=\omega[Tu,v]=Q[u,v].$
Since ${\cal O}$ can be covered by the open sets $\Lambda_{0}(L_{1})$ by varying $L_{1}$ as above (and a fixed $L_{0}$), the above claim proves the continuity of $\lambda$. The second part of the lemma follows now from Lemma 4.1 and the fact that $B_{\rm sym}(L_{0})\rightarrow B_{\rm sym}(V)$, $Q\mapsto Q|_{V}$, preserves Fredholmness of forms since $V$ has finite codimension on $L_{0}$.
4.3    Proof of Theorem 1.2.
Since $\ell([a,b])\subset{\cal O}$, by compactness of $[a,b]$ we can find $a=t_{0}<t_{1}<\cdots<t_{k}=b$ and Lagrangians $L_{1}$, …, $L_{k}$ in $\Lambda_{0}(L_{0})$ such that $\ell([t_{i-1},t_{i}])\subset\Lambda_{0}(L_{i})$ and $L_{i}\supset\mathds{W}^{\perp_{\omega}}$ for all $i$. Fixed $i$, we shall follow the notation in the proof of Lemma 4.3 with $L_{1}$ replaced by $L_{i}$. So, for instance, we have $\mathds{W}/\mathds{W}^{\perp_{\omega}}=\overline{L}_{0}\oplus\overline{L}_{i}$, and it was shown that ¯$\overline{\ell}([t_{i-1},t_{i}])\subset\Lambda_{0}(\overline{L}_{i})$. For each $t\in[t_{i-1},t_{i}]$, let $Q_{t}=\varphi_{L_{0},L_{i}}(\ell(t))\in B_{\rm sym}(L_{0})$ and $\overline{Q}_{t}=\varphi_{\overline{L}_{0},\overline{L}_{i}}(\overline{\ell}(t))\in B_{\rm sym}(V)$. According to the proof of Lemma 4.3, $\overline{Q}_{t}=Q_{t}|_{V}$. Now, from 3. of Theorem 4.2, we have
$\mu_{L_{0}}(\ell,[t_{i-1},t_{i}])-\mu_{\overline{L}_{0}}(\overline{\ell},[t_{i-1},t_{i}])={\rm sf}(Q,[t_{i-1},t_{i}])-{\rm sf}(\overline{Q},[t_{i-1},t_{i}]).$
Since $\overline{Q}_{t}=Q_{t}|_{V}$, the right-hand side above is given by formula (2). Therefore, by summing in $i$, the proof of Theorem 1.2 will be completed once we show that, for $t\in[t_{i-1},t_{i}]$,
$&\ker Q_{t}=\ell(t)\cap L_{0},\\ &V^{\perp_{Q_{t}}}=\pi\bigl{(}\ell(t)\cap(L_{0}+\mathds{W}^{\perp_{\omega}})\bigr{)},\;\;\mbox{hence}\;\;V\cap V^{\perp_{Q_{t}}}=\pi\bigl{(}\ell(t)\cap(V+\mathds{W}^{\perp_{\omega}})\bigr{)},\\ &Q_{t}\big{|}_{V^{\perp_{Q_{t}}}}=\mathfrak{q}_{t}.\end{aligned}\right.$
By writing $\ell(t)$ as the graph of some $T:L_{0}\rightarrow L_{i}$, the proof of the above equalities becomes straightforward and shall be left to the reader.
Data availability statement
No new data were created or analyzed in this study. Data sharing is not applicable to this article.